OpenStudy (anonymous):

I'm working on a practice quiz for Geometry, and I came across two questions that I didn't quite understand. Both follow the same guidelines, so I'll only post the first. "W and D are on circle R. m(length)WRD = 112 degrees. If the area of the sector created by arc WD is 250.21 units squared, what is the diameter of circle R?" I've understood all 38 of the questions other than the two like this. I'm not just asking for an answer, I'd like to know how I can solve this for future reference. Thanks

OpenStudy (anonymous):

Alright, I took Geometry last year and I aced this stuff so if you give me maybe 10 minutes to try to remember and work it out myself first maybe I could help?

OpenStudy (anonymous):

Much appreciated, thanks.

OpenStudy (anonymous):

Was there an illustration that went along with this, or no?

OpenStudy (anonymous):

No illustration. Fairly confusing!

OpenStudy (anonymous):

112/250.21 = 360/x solve for x. x is the total area.

OpenStudy (anonymous):

the angle of the arc is proportional to the area it encloses, for a given circle.

OpenStudy (anonymous):

alright, yeah

OpenStudy (anonymous):

strange thing is that my course actually never covered this.

OpenStudy (anonymous):

geometry is entirely about relationships. ratios. proportionality. congruency. always look for those kinds of relationships, a/b=A/B.

OpenStudy (anonymous):

Well, it's not really uncommon to come across a few things that aren't mentioned in your course. When I took Geometry, we covered the arcs and sectors, the areas and the lengths of these things. I'm trying to pencil a sketch to help me. I still have all of my notebooks they are just at my mother's house. So, if I can't figure this out now, and no one answers your question by the time I go back to her house: I will send you an attachment of some sorts with the notes covering this criteria.

OpenStudy (anonymous):

Thank you very much, both you and Mr_Giblets

OpenStudy (anonymous):

this is my first post here, it's a very nice community :)

OpenStudy (anonymous):

np. geometry can be very difficult. practice doesn't help much, which is very different from calculus and algebra, where you're trying to memorize an algorithm. geometry is best learned through contemplation.

OpenStudy (anonymous):

yeah, I took Alg 2 last year and am still used to, like you said, solving with an algorithm

OpenStudy (anonymous):

Geometry was so much easier to me compared to Alg2 and now I'm in PreCalc and it's just awful! It is very nice! If you ever have any questions involving Geometry I'd be glad to help if I can!

OpenStudy (anonymous):

Thanks a bunch!